# Lifted Forward Planning in Relational Factored Markov Decision Processes with Concurrent Actions

### v2

@@ 7. Empirical Evaluation
Overall, (Approximate) Foreplan achieves a speedup of several orders of magnitude and is able to compute policies for significantly more objects within the same time and memory limits. Moreover, the policy given by Approximate Foreplan is for more than $97$ % of all actions in the epidemic example correct and for all actions in the SysAdmin example.
@@ 8. Conclusion
Propositional planning approaches struggle with having numerous indistinct objects and actions on subsets of those objects. While first-order MDPs can cope with numerous objects, current approaches for solving planning in them neglect concurrent actions and still need to represent actions for each subset individually, resulting in exponentially many actions. In this paper, we present Foreplan, a relational forward planner solving the exponential explosion by lifting the objects: Foreplan groups indistinguishable objects by a representative and carries out calculations on a representative-level. Afterwards, the result is projected to the whole group. Using histograms and focusing only on the number of objects an action is applied to, we effectively reduce the action space from exponential to polynomial in the number of objects.
@@ A.1. Examples for Hyperoperations in Relational Factored MDPs
Definition 3 requires the rewards to be represented as a sum of local reward functions. In fact, this requirement can be lifted to a broader one: Any operation $\oplus$ can run over a collection of local reward functions, as long as this operation has a hyperoperation $\bigoplus$, which performs $\oplus$ repeatedly. The most basic example is addition ($\oplus=+$) and multiplication as the corresponding hyperoperation ($\bigoplus=\cdot$). Moreover, these operations can be linked together in any way: We compute each operation on a collection of local reward functions lifted and then compute the overall result following the linking.
@@ A.6. Example of Removing a Maximum Operator
Let us briefly illustrate the step of removing the maximum operator. The removal is a two-phase process. In a first phase, ALP eliminates the variables and in a second phase, ALP generates the constraints for the linear programm along the elimination sequence. Suppose we have the function
@@ B.4. Instantiation of Linear Program
where $x_{1},x_{2}\in\{0,1,2,3\}$ specify the number of persons being sick and travelling, respectively. The variable $x_{3}$ is Boolean and stores the truth value of $Epidemic$. With $G_{i}$ and $H_{i}$ we denote the lifted computation of $g_{i}$ and $h_{i}$. In the following, we write the lifted computation in terms of $x_{i}$ and $g_{i}$ or $h_{i}$. In the remainder of this subsection, we show the complete constraint generation for the linear program for one example action. We omit all other actions because of limited insights compared to one example instantiation and constraint generation. We show the constraint generation for the action of restricting nobody. We start with writing the constraint tailored to this setting and directly computing the value of basis functions and their backprojections lifted:

### v3

@@ 7. Empirical Evaluation
Overall, (Approximate) Foreplan achieves a speedup of several orders of magnitude and computes policies for significantly more objects within the same time and memory limits. Moreover, the policy given by Approximate Foreplan is correct for SysAdmin and for more than $97$ % of all actions in the epidemic example.
@@ 8. Conclusion
Propositional planning approaches struggle with having numerous indistinct objects and different types of actions applied concurrently. While first-order MDPs can cope with numerous objects, current approaches for solving planning in them neglect concurrent actions and still need to represent actions for each subset individually, resulting in exponentially many actions. In this paper, we present Foreplan, a relational forward planner solving the exponential explosion by lifting the objects: Foreplan groups indistinguishable objects by a representative and carries out calculations on a representative-level. Afterwards, the result is projected to the whole group. Using histograms and focusing only on the number of objects an action is applied to, we effectively reduce the action space from exponential to polynomial in the number of objects.
@@ A.1. Examples for Hyperoperations in Relational Factored MDPs
Definition 3 requires the rewards to be represented as a sum of local reward functions. In fact, this requirement can be lifted to a broader one: Any operation $\oplus$ can run over a collection of local reward functions, as long as this operation has a hyperoperation $\bigoplus$, which performs $\oplus$ repeatedly. The most basic example is addition ($\oplus=+$) and multiplication as the corresponding hyperoperation ($\bigoplus=\cdot$). Moreover, these operations can be linked together in any way: We compute each operation on a collection of local reward functions in a lifted manner and then compute the overall result following the linking.
@@ A.6. Example of Removing a Maximum Operator
Let us briefly illustrate the step of removing the maximum operator. The removal is a two-phase process. In a first phase, ALP eliminates the variables and in a second phase, ALP generates the constraints for the linear program along the elimination sequence. Suppose we have the function
@@ B.4. Instantiation of Linear Program
where $x_{1},x_{2}\in\{0,1,2,3\}$ specify the number of persons being sick and travelling, respectively. The variable $x_{3}$ is Boolean and stores the truth value of $Epidemic$. With $G_{i}$ and $H_{i}$ we denote the lifted computation of $g_{i}$ and $h_{i}$. In the following, we write the lifted computation in terms of $x_{i}$ and $g_{i}$ or $h_{i}$. In the remainder of this subsection, we show the complete constraint generation for the linear program for one example action. We omit all other actions because of limited insights compared to one example instantiation and constraint generation. We show the constraint generation for the action of restricting nobody. We start with writing the constraint tailored to this setting and directly computing the value of basis functions and their backprojections in a lifted way: